OpenStudy (anonymous):

PLEASE HELP ME AND I WILL GIVE MEDAL! Factor v^2-36 A) (v-6)^2 B) (v^2-6)^2 C) (v-6)(v-6) D) (v+6)(v-6) Factor x^2+5x-6 A) (x-1)(x+6) B) (x-6)(x+1) C) (x-3)(x+2) D) (x-2)(x+3)

OpenStudy (anonymous):

the 1st ans uses the formula a^2 - b^2 = (a+b) (a-b)

OpenStudy (anonymous):

can u now tell me wat the ans is??

OpenStudy (anonymous):

I can't sorry bad at this

OpenStudy (solomonzelman):

Use the rule, a² - b² = (a - b)(a + b) v² - 36 = fill in your numbers, your ` a ` is ` v `, AND your ` b ` is ` 6 `.

OpenStudy (anonymous):

Oh wait is it (v+6)(v-6)?

OpenStudy (solomonzelman):

YES good !!!!!!!!!!!

OpenStudy (solomonzelman):

Another one, (imaginary ) v² + 36 = (v - 6i) (v + 6i)

OpenStudy (anonymous):

Thanks

OpenStudy (solomonzelman):

anytime